cheyennelane001:

A florist is preparing a flower arrangement for a wedding. The client wants 6 roses for every 2 carnations in the flower arrangement. If the florist has used 46 carnations in all, how many roses did she use?

dude:

Use a proportion to solve this For every 6 roses there is 2 carnations can be written as \(\large\frac{6}{2}\) Now the question is asking in the place in which the florist has 46 carnations... \(\large[\frac{roses}{carnations}]=>\frac{6}{2}=\frac{42}{?}\) Do you know how to solve this?